OpenStudy (anonymous):

Cards numbered 1 through 30 are placed in a bag and shuffled. One card is chosen at random. What is the probability of randomly selecting a number less than or equal to 5, or a multiple of 7?

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

in the 30 cards, there are only 5 which are less than or equal to 5. so that would be put into the fraction: 1/6 how many multiples of 7 are n 30, do you think?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

7, 14, 21,28

OpenStudy (anonymous):

PLZ HURRY I NEED THIS TURNED IN IN 5 MIN

OpenStudy (anonymous):

so thats 4/30, or 2/15. so i believe you add 2/15 to 1/6. gcf=60. so you have 8/60+ 10/60, or = 18/60. simplify to 3/10

OpenStudy (anonymous):

Thats not one of the answers

OpenStudy (anonymous):

what bout 1/45?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

if this isnt an answer, i give up.. 2/5?

OpenStudy (anonymous):

give up then

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

do u think u could help me on one more?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

.A game at a county fair has a spinning drum holding 15 colored plastic balls: some white, some green, and some red. Every time the drum spins, one ball drops out, its color is recorded, and then the ball is placed back in the drum. The table shows the results after 15, 150, and 1500 spins. What is the best estimate of the number of white balls in the drum? Outcome Number of trials 15 150 1500 white 5 61 680 green 9 74 713 red 1 15 107 A. 1 B. 4 C. 6 D. 7

OpenStudy (anonymous):

wait for the first one, i think its 3/10

OpenStudy (anonymous):

its not one of the answers though

OpenStudy (anonymous):

can u answer the second one?